The nearest star To our sun is Proxima Centauri, at a distance of 4.3 light years from the sun. A light year is the distance that light travels in one year. How far away in kilometers is Próxima centaury from the sun?
What formula do you use and step by step solution
1
Expert's answer
2020-09-29T06:12:07-0400

The nearest star To our sun is Proxima Centauri, at a distance of 4.3 light years from the sun. A light year is the distance that light travels in one year. How far away in kilometers is Próxima centaury from the sun?

What formula do you use and step by step solution?

Solution:

As we know 1 light year equals 9.461*10^12 kilometers.

So make a proportion:

1 light year - 9.461*10^12 km

4.3 light years - X km

Where:

X = 4.3*9.461*1012/1 = 40.6823*10^12 km

That’s why Próxima centaury is from the sun at a distance 40.6823*10^12 kilometers.

Answer: Próxima centaury